Maria left the starting line at 8:30 a.m. and drove at an average rate of 40 mi/h. What time did she arrive at Checkpoint 1? (the entire route from the start line to the finish line is a 24 mile circle)
Answer by stanbon(75887)

Maria left the starting line at 8:30 a.m. and drove at an average rate of 40 mi/h. What time did she arrive at Checkpoint 1? (the entire route from the start line to the finish line is a 24 mile circle)
----------------------------
time = distance/rate = 24/40 = 3/5 hr = (3/5)(60 min) = 36 minutes
-----------------
Arrival time = 8:30 + 36 min = 9:06 AM
========================================
Cheers,
Stan H.